# A NEW LOWER BOUND FOR THE DIAGONAL POSET RAMSEY NUMBERS

### v1

@@ 1 Introduction
$R(Q_{n},Q_{n})\leq n^{2}-\left(1-\frac{2}{\sqrt{\log n}}\right)\frac{n}{\log n}.$
@@
For $n$ large enough, there exists a constant $k$ such that $R(Q_{n},Q_{n})>2.7n+k$.
@@
Our proof is inspired by Winter’s proof in which a $4$-layered colouring is modified with the help of two families of sets, called pivots. Our strategy is to start with a large number of layers, pair them up, and inside each pair modify the colouring according to two families of pivots tailored for said pair. The pivots (and their parameters) are designed to force a monochromatic embedding to ‘skip’ a certain number of levels. If one manages to force a monochromatic embedding to skip more than $n+1$ levels in total, there is, of course, no monochromatic $Q_{n}$.
In Winter’s proof, the existence of pivots relies exclusively on probabilistic methods. Our proof is more constructive in the sense that we pinpoint exactly one type of families of pivots. This gives more control and, in consequently, tighter bounds. It turns out that starting with 602 layers is enough to achieve the claimed $2.7n$ bound. Moreover, for an arbitrary number of initial layers, we pinpoint the exact restrictions that need to be satisfied in order for the proof to go through, which suggest that the more layers, the better the bound, although not exceeding $3n$.
@@ 2 A lower bound of $(2+\frac{1}{3})n$
Since $\phi$ is an embedding, we then get $\phi(\{i\})\subseteq\phi([n]\setminus\{i\})$, a contradiction. Therefore, for every $i\in[n]$ there exists $a(i)\in\phi(\{i\})\setminus\bigcup_{A\subseteq([n]\setminus\{i\})}\phi(A)$. We further notice that if $i\neq j$ $\phi(\{j\})$ and $\bigcup_{A\subseteq([n]\setminus\{i\})}\phi(A)$ are disjoint as $\{j\}\subseteq[n]\setminus\{i\}$. As such, $a(i)\neq a(j)$ if $i\neq j$. Let $X=\{a(i):i\in[n]\}$, which is in bijection with $[n]$. Given $A\subseteq X$, let $\hat{A}=\{a^{-1}(x):x\in A\}$. We therefore define $\phi^{\prime}:\mathcal{Q}(X)\to Q$ as follows
@@ 3 Existence of pivot sets
Together with the fact that $|\mathcal{K}^{\prime}_{s}|\geq(1.0002)^{n}$, we have that $\mathbb{P}[\mathcal{K}_{s}\text{ is bad}]\leq(0.9997)^{n(1.0002)^{n}}\to 0$ as $n\to\infty$, which finishes the claim. ∎
@@ 4 Proof of the main result
For $n$ large enough, there exists a constant $k$ such that $R(Q_{n},Q_{n})>2.7n+k$.
@@ 5 Limitations and future directions
Let $\epsilon>0$. Then there exists $L$ and a constant $a$ such that the above type of colouring ($2L$ pairs of layers modified by 2 families of pivots) gives, for large enough $n$,
@@
In this appendix we explain the Python code used to optimise the parameters that give us Theorem 1. The parameters we optimise over are a $c_{i}$ value and a $h_{i}$ value for each layer. The code sets up the relevant constraints for these parameters as discussed in the proof outline of Theorem 1, and then calls a function from SciPy to maximise the total value of $c=2\sum_{i=1}^{L}c_{i}$.
@@
14# bottom = bottom level of the layer as frac of n, doesn’t include contributions from c
15# top = top level of layer as frac of n, doesn’t include contributions from c
16# red_level = what frac of n have we already climbed in previous layers for red at the bottom
17# red_climb = what frac of n are we allowed to climb in the red part for this layer
18# blue_level = what frac of n have we already climbed in previous layers for blue at the top
19# blue_climb = what frac of n are we allowed to climb in the blue part for this layer
20Layer = namedtuple("Layer", ["index", "bottom", "top", "red_level", "red_climb", "blue_level", "blue_climb", "num_layers"])
21
22def make_layers(num_layers):
@@
37
38# top = bottom + red_climb + blue_climb
39# = red_level + blue_level + red_climb
40# bottom = red_level + blue_level - blue_climb
@@
41# Define N
42def var_N(vars, layer: Layer):
@@
45
46# Define s
47def var_s(vars, layer: Layer):
@@
50
51# Define t
52def var_t(vars, layer: Layer):
@@
56
57# Define the top level of a layer including contributions from c
58def var_top(vars, layer: Layer):
@@
61
62# Calculate K_s
63def Ks_log(vars, layer: Layer):
@@
65
66# Calculate K_t
67def Kt_log(vars, layer: Layer):
@@
71
72# Calculate t-neighbourhood of an s-set
73def Nt_log(vars, layer: Layer):
@@
75
76# Calculate how many sets in K_s have union of size at most t with a specified set.
77def Nsect_log(vars, layer: Layer):
@@
82# We need level t to be inside the layer, so t <= top -> top - t >= 0
83# This constraint should be covered by the Nsect_max_constraint, so
84# we don’t include it in the optimisation, but we can check it afterwards
85def t_lessThan_Top_constraint(vars, layer: Layer):
@@
88
89# I-Constraint: This comes from finding a large subset of K_s(P,X) with on large intersections
90# h+c <= RC - RC^2/(1-RL) -> RC - RC^2/(1-RL) - (h+c) >= 0
91def Nsect_max_constraint(vars, layer: Layer):
@@
95
96# p-constraint: This combines the constraints that 1 <= pt * Kt and pt * Nt <= 1,
97# which is equivalent to Kt/Nt >= 1 -> Kt - Nt >= 0
98def room_for_pt_constraint(vars, layer: Layer):
@@
100
101# Constraint: 1 <= K’ -> K’ - 1 >= 0
102# This constraint should be covered by the Nsect_max_constraint, so
103# we don’t include it in the optimisation, but we can check it afterwards
@@
[…]
126
127Finally, we setup our objective function (to maximise $\sum$_i c_i\$) and run the optimiser in a loop while ensuring all constraints hold.
128
[…]
130# Objective function: maximise c (equivalent to minimising -c)
131def objective(vars, num_layers):
@@
215 return best, best_vars
216
217best, best_vars = optimiseForNLayers(150)
218
219print("#############################################")
220print("Best: ", best)
221print("Best parameters: ", best_vars)

### v2

@@ 1 Introduction
$R(Q_{n},Q_{n})\leq n^{2}-\left(1-\frac{2}{\sqrt{\log n}}\right)n\log n.$
@@
There exists a constant $k$ such that $R(Q_{n},Q_{n})>2.7n+k$.
@@
Our proof is inspired by Winter’s proof in which a $4$-layered colouring is modified with the help of two families of sets, called pivots. Our strategy is to start with a large number of layers, pair them up, and inside each pair modify the colouring according to two families of pivots tailored for said pair. The pivots (and their parameters) are designed to force a monochromatic embedding to ‘skip’ a certain number of levels. If one manages to force a monochromatic embedding to skip enough levels in total, there is, of course, not enough levels left to embed a monochromatic $Q_{n}$.
In Winter’s proof, the existence of pivots relies exclusively on probabilistic methods. Our proof is more constructive in the sense that we pinpoint exactly one type of families of pivots. This gives more control and consequently tighter bounds. It turns out that starting with 602 layers is enough to achieve the claimed $2.7n$ bound. Moreover, for an arbitrary number of initial layers, we pinpoint the exact restrictions that need to be satisfied in order for the proof to go through, which suggests that the more layers, the better the bound, although not exceeding $3n$.
@@ 2 A lower bound of $(2+\frac{1}{3})n$
Since $\phi$ is an embedding, we then get $\phi(\{i\})\subseteq\phi([n]\setminus\{i\})$, a contradiction. Therefore, for every $i\in[n]$ there exists $a(i)\in\phi(\{i\})\setminus\bigcup_{A\subseteq([n]\setminus\{i\})}\phi(A)$. We further notice that if $i\neq j$, $\phi(\{j\})$ and $\phi(\{i\})\setminus\bigcup_{A\subseteq([n]\setminus\{i\})}\phi(A)$ are disjoint as $\{j\}\subseteq[n]\setminus\{i\}$. As such, $a(i)\neq a(j)$ if $i\neq j$. Let $X=\{a(i):i\in[n]\}$, which is in bijection with $[n]$. Given $A\subseteq X$, let $\hat{A}=\{a^{-1}(x):x\in A\}$. We therefore define $\phi^{\prime}:\mathcal{Q}(X)\to Q$ as follows
@@ 3 Existence of pivot sets
Together with the fact that $|\mathcal{K}^{\prime}_{s}|\geq(1.0002)^{n}$, we have that $\mathbb{P}[\mathcal{K}_{s}\text{ is bad}]\leq(0.9997)^{n(1.0002)^{n}}$, which finishes the claim. ∎
@@ 4 Proof of the main result
There exists a constant $k$ such that $R(Q_{n},Q_{n})>2.7n+k$.
@@ 5 Limitations and future directions
Let $\epsilon>0$. Then there exists $L$ and a constant $a$ such that the above type of colouring ($2L$ pairs of layers modified by 2 families of pivots) gives
@@
In this appendix we explain the Python code used to optimise the parameters that give us Theorem 1. This code can also be found at this GitHub repository. The parameters we optimise over are a $c_{i}$ value and a $h_{i}$ value for each layer. The code sets up the relevant constraints for these parameters as discussed in the proof outline of Theorem 1, and then calls a function from SciPy to maximise the total value of $c=2\sum_{i=1}^{L}c_{i}$.
@@
15# bottom = bottom level of the layer as frac of n, doesn’t include contributions from c
16# top = top level of layer as frac of n, doesn’t include contributions from c
17# red_level = what frac of n have we already climbed in previous layers for red at the bottom
18# red_climb = what frac of n are we allowed to climb in the red part for this layer
19# blue_level = what frac of n have we already climbed in previous layers for blue at the top
20# blue_climb = what frac of n are we allowed to climb in the blue part for this layer
21Layer = namedtuple("Layer", ["index", "bottom", "top", "red_level", "red_climb", "blue_level", "blue_climb", "num_layers"])
22
23def make_layers(num_layers):
@@
38
39# top = bottom + red_climb + blue_climb
40# = red_level + blue_level + red_climb
41# bottom = red_level + blue_level - blue_climb
@@
43# Define N
44def var_N(vars, layer: Layer):
@@
47
48# Define s
49def var_s(vars, layer: Layer):
@@
52
53# Define t
54def var_t(vars, layer: Layer):
@@
58
59# Define the top level of a layer including contributions from c
60def var_top(vars, layer: Layer):
@@
63
64# Calculate K_s
65def Ks_log(vars, layer: Layer):
@@
67
68# Calculate K_t
69def Kt_log(vars, layer: Layer):
@@
73
74# Calculate t-neighbourhood of an s-set
75def Nt_log(vars, layer: Layer):
@@
77
78# Calculate how many sets in K_s have union of size at most t with a specified set.
79def Nsect_log(vars, layer: Layer):
@@
85# We need level t to be inside the layer, so t <= top -> top - t >= 0
86# This constraint should be covered by the Nsect_max_constraint, so we don’t include it in the optimisation, but we can check it afterwards
87def t_lessThan_Top_constraint(vars, layer: Layer):
@@
90
91# I-Constraint: This comes from finding a large subset of K_s(P,X) with on large intersections
92# h+c <= RC - RC^2/(1-RL) -> RC - RC^2/(1-RL) - (h+c) >= 0
93def Nsect_max_constraint(vars, layer: Layer):
@@
97
98# p-constraint: This combines the constraints that 1 <= pt * Kt and pt * Nt <= 1, which is equivalent to Kt/Nt >= 1 -> Kt - Nt >= 0
99def room_for_pt_constraint(vars, layer: Layer):
@@
101
102# Constraint: 1 <= K’ -> K’ - 1 >= 0
103# This constraint should be covered by the Nsect_max_constraint, so we don’t include it in the optimisation, but we can check it afterwards
@@
Finally, we setup our objective function (to maximise $\sum_{i}c_{i}$) and run the optimiser in a loop while ensuring all constraints hold.
126# Objective function: maximise c (equivalent to minimising -c)
127def objective(vars, num_layers):
@@
215print("#############################################")
216print("Best: ", best)
217print("Best parameters: ", best_vars)